Last question of the day. I need to check my work on the following equation:
0.03(3-7x)-0.02(1-x)=0.64
I solved the equation which is -3, but I cannot for the life of me figure out how to check this answer. I have checked my textbook, and see no examples. I know I should place -3 in place of x, but it doesn't seem to work.
I have tried 0.03(3-7(-3))-0.02(1-(-3))=0.64 and 0.03(3-7-3)-0.02(1-3)=0.64
each do not produce a TRUE answer.
Thank you for your help.

Answer by ikleyn(53937)

.
To check your answer, simply substitute the value x= -3 into the expression. You will get


0.03*(3-7*(-3)) - 0.02*(1-(-3)) = 0.03*(3+21) - 0.02*(1+3) = 0.03*24 - 0.02*4 = 0.72 - 0.08 = 0.64.


How you do present your calculations, shows that you have serious problem with arithmetic.


Simply saying, you don't know the basic arithmetic rules.



-----------------
Comment from student:  Yes,  this is my first math class in 20 years.  I’m learning what I didn’t in High School.  Thank you for the clarification.
-----------------


My response:    O-o-o !    I understand.

Then come to this forum with your questions/problems often !

We will be glad to help you !


                    H a p p y   l e a r n i n g  ! !